# An Exploration of Default Images in Text-to-Image Generation

### v2

@@ 1. Introduction
We argue that studying default images is important for the creative practice of prompt engineering for TTI. When a generative model responds with default images, it is a clear indication that the prompt given to the model is not understood by the model. At the same time, default images reveal gaps in the TTI model’s vocabulary and limits in its capabilities. Consider, for instance, a creative practitioner prompting a generative model with the Tagalog word “pagpag” (referring to a dish made from leftover food scavenged from garbage dumps) and receiving a default image back (e.g., a painting of a woman with a cloud above her face). This is a clear indication that term was not understood and the practitioner should revise the wording of the prompt to improve the generated images. Default images can put an early end to a potentially unfruitful pathway in the practitioner’s prompting of the model. However, default images could also derail the creative process and the quest to generate the desired image. Therefore, default images and their recognition play a crucial role in the creative practice of prompt engineering and for discovering the limits of what a TTI model is capable of. To date, default images have not been investigated in prior research.
In this paper, we present the first preliminary exploration into default images in Midjourney111www.midjourney.com, a popular image generation system. We describe our approach to systematically create a set of input prompts likely triggering default images, and describe our experiments in generating images from these prompts. We then apply affinity diagramming to identify similarities between the generated images, resulting in a set of ten default images. We describe these images and make nine postulates about their characteristics. We further conducted a user study investigating whether default images affect user satisfaction, and how they align with human perception and expectations. Note that due to the vastness of Midjourney’s latent space, this initial investigation is not exhaustive. We highlight the challenges of identifying unique but realistic input prompts that reliably trigger default images in Midjourney, and the importance of studying default images.
@@
  • A crowdsourced user study investigating how default images affect user satisfaction.
@@
2. Importance of Studying Default Images in Text-to-Image Generation
@@ 3.3. Experiments and User Studies on Generated Image Quality and Satisfaction
4. Default Images in Text-to-Image Generation
4.1. What are Default Images?
@@ 5.1. Input Prompts Likely Triggering Default Images
We then turned to create prompts that combine both not recognized terms and practical use cases, which we deem most promising to our endeavor. These prompts are likely used in the real world, but still challenge the model. This includes:
@@
The six categories of prompts emphasize real-world usefulness and they are our focus area and inspiration for creating default images. The low-resource languages were chosen for their language-native words that are not borrowed from other languages and, thus, do not cognate with widely known foreign languages. We focus on Finnish in this study (see Appendix A4.1), leaving Tagalog (Appendix A4.2) to future work. Using this approach, we created a set of prompts, listed in Appendix A. Each of the prompts was checked on haveibeentrained.com, a website that allows to search the training data used in many TTI models (including LAION (LAION.ai, nd; Schuhmann et al., 2022)). Prompts were only included if a search did not yield large numbers of images as a search result. We then generated images with these prompts on Midjourney, as described in the following section.
@@ 6.3.1. Participants
The survey participants ($N=30$) were between 21 and 53 years old (Mean=30.1, SD=6.8) and included 17 men and 13 women. About 53% of the participants had completed a Bachelor degree and 23% a Master’s degree. Participants were familiar with TTI ($Mean=5.8$ and $SD=1.2$ on a scale of 1 (Strongly Disagree) to 7 (Strongly Agree).
@@ 6.3.2. User study results
Figure 6 shows that unknown terms potentially triggering default images may negatively affect user satisfaction, but not in all cases. In Q1, the effect of replacing ‘raven’ with ‘greagoft’ resulted in only subtle changes to the resulting image. The image appeared almost as if a raven could be there. The participants’ satisfaction, therefore, was high ($Mean=5.2$, $SD=1.8$). In Q2, where ‘leaves and flowers’ was replaced with ‘greagoft’, the divergence between prompt and resulting image was more noticeable and satisfaction was lower ($Mean=2.7$, $SD=1.9$). A Fisher’s exact test indicates that the way participants perceived the change (subtle vs. noticeable) is significantly related to their satisfaction level ($OR=0.074$, $p<0.05$). The odds ratio of 0.074 indicates that dissatisfaction (1–3 on the Likert scale) is much less likely when the change is subtle compared to noticeable. In other words, participants were significantly more likely to be satisfied (5–7) when the change in the image was subtle (Q1) rather than noticeable (Q2).
For the third question (Q3), we did not switch or omit any words in the prompt, and the image was the exact result of the prompt. Participants were satisfied with the resulting image ($Mean=4.5$, $SD=1.9$). In the fourth question (Q4), we used a prompt known to trigger default images. Participants showed particularly strong dissatisfaction with this image ($Mean=2.6$, $SD=2.0$). Again, there was a significant difference in user satisfaction between Q3 (representing an exact match between prompt and image) and Q4 (respresenting a default image) with $OR=0.21$ and $p<0.05$.
Figure 6. Survey responses, showing how default images can lead to user dissatisfaction.
@@ 7.3. Explainability and Alignment of Text-to-Image Generation
Because it is often not clear why the model produced a certain default image, the problem of default images touches on research in the field of interpretability and Explainable AI (XAI). In a sense, the default image – currently – is the mechanism of the TTI system to let the user know that the prompt is not understood. Future work could investigate how a user should be alerted about such prompts, e.g. by designing user interfaces that provide feedback to the user about which parts of the prompt was not understood. Such a solution could help to write better prompts, reducing the appearance of the default images, and thus saving time in the process of prompt engineering. The reduction of both the real and the perceived occurrence of default images would help in producing results that are more satisfying to human perception.
7.4. Limitations and Future Work

### v3

@@ 1. Introduction
We argue that studying default images is important for the creative practice of prompt engineering for TTI. Default images reveal gaps in the TTI model’s vocabulary and capabilities. Consider, for instance, a creative practitioner prompting a generative model with the Tagalog word “pagpag” (referring to a dish made from leftover food scavenged from garbage dumps) and receiving a default image back (e.g., a painting of a woman with a cloud above her face). This is a clear indication that term was not understood and the practitioner should revise the wording of the prompt to improve the generated images. Default images can put an early end to a potentially unfruitful pathway in the practitioner’s prompting of the model. However, default images could also derail the creative process and the quest to generate the desired image. Therefore, default images and their recognition play a crucial role in the creative practice of prompt engineering and for discovering the limits of what a TTI model is capable of. To date, default images have not been investigated in prior research.
In this paper, we present the first exploration into default images in Midjourney111www.midjourney.com, a popular image generation system. We describe our approach to systematically create a set of input prompts likely triggering default images, and describe our experiments in generating images from these prompts. We then apply affinity diagramming to identify similarities between the generated images, resulting in a set of ten default images. We describe these images and make nine postulates about their characteristics. We further conducted a user study investigating whether default images affect user satisfaction, and how they align with human perception and expectations. Note that due to the vastness of Midjourney’s latent space, this initial investigation is not exhaustive. We highlight the challenges of identifying unique but realistic input prompts that reliably trigger default images in Midjourney, and the importance of studying default images.
@@
  • A user study with crowdsourced participants ($N=48$) investigating how default images affect user satisfaction.
@@
2. Default Images in Text-to-Image Generation
2.1. What are Default Images?
2.2. Importance of Studying Default Images in Text-to-Image Generation
@@ 4.1. Input Prompts Likely Triggering Default Images
We then turned to create prompts that combine both unrecognized terms and practical use cases, which we deem most promising to our endeavor. These prompts are likely used in the real world, but still challenge the model. This includes:
@@
The six categories of prompts emphasize real-world usefulness and they are our focus area and inspiration for creating default images. The low-resource languages were chosen for their language-native words that are not borrowed from other languages and, thus, do not cognate with widely known foreign languages. We focus on Finnish in this study (see Appendix A4.1), leaving Tagalog (Appendix A4.2) to future work. Using this approach, we created a set of prompts, listed in Appendix A. To assess the likelihood that our prompts had been encountered during model training, we checked each prompt on haveibeentrained.com, a search interface for over 5.8 billion image-caption pairs. This tool enables users to determine whether specific terms appear in large-scale datasets, such as LAION (LAION.ai, nd). These datasets are commonly used in the training of text-to-image models (LAION.ai, nd; Schuhmann et al., 2022). We used this website to identify prompts that were absent from or underrepresented in the training data, increasing the likelihood that any generated imagery would result in default or fallback behavior. Our inclusion criteria were twofold: 1) the prompt should plausibly occur in real-world use, and 2) it should pose a challenge to the model due to low training data coverage. Prompts that did not return matches in haveibeentrained.com were prioritized, as these were most likely to elicit unexpected or default imagery. We then generated images with these prompts on Midjourney, as described in the following section.
@@ 5.3.1. Participants
The survey participants ($N=48$) were between 21 and 68 years old (Mean=32.9, SD=9.7) and included 27 men and 21 women. About 52% of the participants had completed a Bachelor degree and 23% a Master’s degree. Participants were familiar with TTI ($Mean=6.0$ and $SD=1.3$ on a scale of 1 (Strongly Disagree) to 7 (Strongly Agree).
@@ 5.3.2. User study results
Figure 6 shows that unknown terms potentially triggering default images may negatively affect user satisfaction, but not in all cases. In Q1, the effect of replacing ‘raven’ with ‘greagoft’ resulted in only subtle changes to the resulting image. The image appeared almost as if a raven could be there. The participants’ satisfaction, therefore, was high ($Mean=4.9$, $SD=2.0$). In Q2, where ‘leaves and flowers’ was replaced with ‘greagoft’, the divergence between prompt and resulting image was more noticeable and satisfaction was lower ($Mean=2.6$, $SD=1.8$). A Fisher’s test with continuity correction indicates that the way participants perceived the change (subtle vs. noticeable) is significantly related to their satisfaction level ($OR=10.7$, $p<0.05$). The odds ratio of 10.7 indicates that dissatisfaction (1–3 on the Likert scale) is much less likely when the change is subtle compared to noticeable. In other words, participants were significantly more likely to be satisfied (5–7) when the change in the image was subtle (Q1) rather than noticeable (Q2).
For the third question (Q3), we did not switch or omit any words in the prompt, and the image was the exact result of the prompt. Participants were satisfied with the resulting image ($Mean=4.6$, $SD=1.8$). In the fourth question (Q4), we used a prompt known to trigger default images. Participants showed particularly strong dissatisfaction with this image ($Mean=2.4$, $SD=1.9$). Again, there was a significant difference in user satisfaction between Q3 (representing an exact match between prompt and image) and Q4 (respresenting a default image) with $OR=16.2$ and $p<0.05$.
Figure 6. Responses from the user study ($N=48$) show how default images can lead to user dissatisfaction.
@@ 6.3. Explainability and Alignment of Text-to-Image Generation
Because it is often not clear why the model produced a certain default image, the problem of default images touches on research in the field of interpretability and Explainable AI (XAI). In a sense, the default image – currently – is the only mechanism of the TTI system to let the user know that the prompt is not understood. Other generative models, such as language models, include clear, explicit ways of letting users know that something is not understood or that they cannot respond to it (e.g., refusing to answer for reasons of content policy). TTI models do not have such mechanisms natively. In practice, language models handle such sanity checks on Midjourney. Our prompts, however, pass these checks which results in the observed default images. Default images obscure the fact that the model can, in fact, not respond to the user prompt as expected. Instead of returning nothing, the TTI model generates default images. This could be termed a case of ostensive inferential communication (Sperber and Wilson, 1995; Wilson and Sperber, 2006), a model of communication where a speaker provides cues (ostension) to make their intended meaning clear, and the listener infers that meaning through a process of reasoning (inference). This is also a plausible explanation as to why user satisfaction is lower when the image is different from expectations.
6.4. Future Work and Limitations
Future work could investigate how a user should be alerted about prompts that are not understood by the TTI model, e.g. by designing user interfaces that provide feedback to the user about which parts of the prompt was not understood. Such a solution could help to write better prompts, reducing the appearance of the default images, and thus saving time in the process of prompt engineering. The reduction of both the real and the perceived occurrence of default images would help in producing results that are more satisfying to human perception.